# Detecting Hyperons in neutron stars – a machine learning approach

### v1

@@ I Introduction
From the theoretical side, the low-density region of the NS EOS is constrained by chiral effective field theory (cEFT) Hebeler et al. (2013); Drischler et al. (2019); Huth et al. (2022), while perturbative quantum chromodynamics (pQCD) becomes reliable at high densities Kurkela et al. (2010) (see Ghiglieri et al. (2020) for a review). Inference of the NS EOS, given a set of theoretical, experimental and observational constraints, is frequently carried out using Bayesian inference frameworks Landry and Essick (2019); Traversi and Char (2020); Raaijmakers et al. (2021); Essick et al. (2021); Huth et al. (2022); Malik et al. (2022); Mondal and Gulminelli (2023); Gorda et al. (2023); Sun et al. (2023); Huang et al. (2023); Malik et al. (2023). Several studies take an agnostic modeling of the EOS, e.g., piecewise polytropes, the speed of sound formalism, spectral approaches, Gaussian processes Kurkela et al. (2014); Annala et al. (2018); Landry and Essick (2019); Tews et al. (2018); Annala et al. (2020); Essick et al. (2020); Somasundaram et al. (2023); Altiparmak et al. (2022); Gorda et al. (2023). Although these methods provide valuable insights into the EOS, they do not offer information about the underlying particle composition of NS matter.
@@ II.1 Theoretical concepts of dataset
The data used to train the machine learning models is derived from nuclear models based on an RMF theory, where nucleon interactions are mediated by the exchange of scalar-isoscalar, vector-isoscalar, and vector-isovector mesons Mueller and Serot (1996); Horowitz and Piekarewicz (2001); Todd-Rutel and Piekarewicz (2005); Providência et al. (2014); Chen and Piekarewicz (2014); Tolos et al. (2017). To allow a realistic description of nuclear matter, the Lagrangian densities include self-interacting and mixed meson terms. These terms define the density dependence of the EOS and the symmetry energy. We refer to this approach and the corresponding EOS as nonlinear (NL). An alternative would be to consider density dependent meson couplings Typel and Wolter (1999). In Sec. IV.3 a set of EOSs obtained within this last formalism in Malik et al. (2022) will also be introduced and referred to as DDB EOSs. A detailed description of the EOS used in the present study can be found in Malik et al. (2023), where within a Bayesian inference calculation the parameters of the RMF model were constrained. The data set, which is publicly available [^1] https://zenodo.org/records/7854112, contains 17810 EoS of pure nucleonic matter and 18728 EoS of hadronic matter containing both nucleons and hyperons. The $\Lambda$ and $\Xi$ hyperon-meson couplings are allowed to vary in a range compatible with the properties of hypernuclei as determined in the studies Fortin et al. (2017); Providência et al. (2018); Fortin et al. (2020). The corresponding mass-radius curves, calculated using the Tolman-Oppenheimer-Volkoff (TOV) equations, are shown in Fig. 1. The Bayesian framework used to generate the two sets of EOSs imposes minimal constraints on various nuclear saturation properties. This approach ensures the reproduction of $2M_{\odot}$ NSs and maintains consistency with low-density pure neutron matter as predicted by N3LO calculations in chiral effective field theory (see table 1).
@@
Figure 1: Mass-radius, left plot, and mass-log(tidal deformability), right plot, curves for nucleonic matter (17810 curves, shown in blue) and hyperonic matter (18728 curves, shown in red).
@@ V Conclusions
We investigated a classification problem aimed at determining the potential presence of hyperons within NSs based on their macroscopic properties. We investigate the capacity of NNs for the present binary classification problem of detecting the presence of hyperons. The dataset used has been constructed from an RMF approach within a Bayesian framework, incorporating crucial constraints derived from both nuclear matter properties and NS observations. The dataset consists of 17810 EOSs for pure nucleonic matter and 18728 EOSs including hyperons. This approach was chosen to achieve our primary objective: to investigate the presence of hyperons within NSs based on observational data, which requires the use of microscopic models. From this comprehensive dataset, we generated three distinct sets by selecting either 5, 10, or 15 pairs of mass-radius ($\bm{R}$), mass-tidal deformability ($\bm{\Lambda}$), or both ($\bm{R\Lambda}$) as input parameters. We varied the input noise to create three different datasets for each input size, as summarized in Table II. These sets represent different levels of noise scatter in the input data. Consequently, a total of 27 NN models were trained, each corresponding to one of these possible combinations. We conducted a grid search to optimize the model architecture specifically for the $R_{0}$, $\Lambda_{0}$, and $R\Lambda_{0}$ sets. Once the optimal architecture was identified, we evaluated the model’s performance using confusion matrices and key metrics such as accuracy and F1 score. To ensure that the model’s performance is not biased by a fixed threshold, we also calculated the AUC for accuracy as a function of the threshold. The AUC values demonstrated consistent behavior with the accuracy measured at the standard threshold of $0.5$ Our findings reveal that as input noise increases, the model requires a larger input size to maintain high accuracy. In contrast, for datasets without noise, the model quickly reaches peak accuracy with a smaller number of input pairs. This suggests that for observational data—roughly equivalent to the sets with $\mathbb{X}_{2}$ noise levels—larger input sizes are needed to achieve accuracy comparable to the noise-free case. Nevertheless, our model achieves high scores across all evaluated metrics, including in scenarios with the highest input noise. We further tested the model on a different dataset, also generated using an RMF approach but with density-dependent couplings and a different particle composition. The results demonstrated good accuracy across all models, consistently exceeding 73%. For the AUC value and the F1 score, the performance decreased slightly. The F1 score is particularly significant given the highly imbalanced nature of this test set, and has indicated that the model is effectively handling and correctly predicting outcomes despite the data imbalance. While the model performs well across most datasets, there is room for improvement, particularly in handling the $\Lambda$ datasets with higher noise levels. Finally, we tested our models on real observational data for the $\bm{R}$ models with $Q=5$, using a subset of observations that lie within the model’s training region. The results revealed that the majority of predictions fell within the region without hyperons, suggesting that hyperons may be absent inside NSs. More specifically, 65% of the data across all datasets indicated predictions in the region without hyperons. We also conducted tests with fewer samples for the $R_{2}$ set at $Q=10$ and $Q=15$, achieving solid results that consistently indicate the absence of hyperons.
@@
Looking ahead, future observations will enable us to test models that incorporate tidal deformability, potentially providing additional insights into NS compositions. A significant direction for future work involves the application of deep ensembles or the incorporation of BNNs. These approaches would allow us to quantify both epistemic and aleatoric uncertainties, providing a more comprehensive understanding of model reliability. Given the results obtained with different nuclear models it could also be advantageous to perform a more rigorous model optimization, such as using Bayesian search methods, to enhance performance and accuracy. However, this approach can be resource-intensive in terms of both time and memory. Additionally, expanding the model to consider the presence of a quark phase in NS compositions and training and testing it under these new conditions could offer intriguing insights into the physics of dense matter.

### v2

@@ I Introduction
From the theoretical side, the low-density region of the NS EOS is constrained by chiral effective field theory (cEFT) Hebeler et al. (2013); Drischler et al. (2019); Huth et al. (2022), while perturbative quantum chromodynamics (pQCD) becomes reliable at high densities Kurkela et al. (2010) (see Ghiglieri et al. (2020) for a review). Inference of the NS EOS, given a set of theoretical, experimental and observational constraints, is frequently carried out using Bayesian inference frameworks Landry and Essick (2019); Traversi and Char (2020); Raaijmakers et al. (2021); Essick et al. (2021); Huth et al. (2022); Malik et al. (2022); Mondal and Gulminelli (2023); Gorda et al. (2023); Sun et al. (2023a); Huang et al. (2023); Malik et al. (2023). Several studies take an agnostic modeling of the EOS, e.g., piecewise polytropes, the speed of sound formalism, spectral approaches, Gaussian processes Kurkela et al. (2014); Annala et al. (2018); Landry and Essick (2019); Tews et al. (2018); Annala et al. (2020); Essick et al. (2020); Somasundaram et al. (2023); Altiparmak et al. (2022); Gorda et al. (2023). Although these methods provide valuable insights into the EOS, they do not offer information about the underlying particle composition of NS matter.
@@ II.1 Theoretical concepts of dataset
In the present study we adopt the description of baryonic matter within a RMF approach Glendenning (1996), a framework which allows to include hyperonic degrees of freedom in a straightforward way. The EOS is obtained from a Lagrangian density written in terms of the Dirac spinors ${\Psi}$ that describe nucleons, and hyperons in the case of hyperonic matter, as well as leptons to describe $\beta$-equilibrium and electrically neutral matter. The baryonic interaction is described in terms of an exchange of mesons: a scalar isoscalar $\sigma$ field, a vector isoscalar $\omega$ field, a vector isovector $\varrho$ field, and in the case of hyperonic matter a vector isoscalar $\phi$ field with hidden strangeness. The meson masses are $m_{i},\,i=\sigma,\,\omega,\,\varrho,\,\phi$. The parameters $g_{i}$, $i=\sigma,\,\omega,\,\varrho,\,\phi$ designate the couplings of the mesons to the baryons. The dataset used in the present study derives from the Lagrangian density
$\mathcal{L}=$ $\bar{\Psi}\Big{[}\gamma^{\mu}\left(i\partial_{\mu}-g_{\omega}A_{\mu}^{(\omega)}-g_{\varrho}{\bm{t}}\cdot\bm{A}_{\mu}^{(\varrho)}\right)-\left(m-g_{\sigma}\phi\right)\Big{]}\Psi$ (1)
$+\frac{1}{2}\left(\partial_{\mu}\phi\partial^{\mu}\phi-m_{\sigma}^{2}\phi^{2}\right)$
$-\frac{1}{4}F_{\mu\nu}^{(\omega)}F^{(\omega)\mu\nu}+\frac{1}{2}m_{\omega}^{2}A_{\mu}^{(\omega)}A^{(\omega)\mu}$
$-\frac{1}{4}\bm{F}_{\mu\nu}^{(\varrho)}\cdot\bm{F}^{(\varrho)\mu\nu}+\frac{1}{2}m_{\varrho}^{2}\bm{A}_{\mu}^{(\varrho)}\cdot\bm{A}^{(\varrho)\mu}+\mathcal{L}_{NL}.$
The last term $\mathcal{L}_{NL}$ includes all the non-linear (NL) mesonic terms and is absent in frameworks with density dependent couplings $g_{i}$. Our dataset is generated considering this term in the Lagrangian density together with constant $g_{i}$ meson-baryon couplings, this approach and the corresponding EOS are referred to as NL. An alternative would be to consider density dependent meson couplings Typel and Wolter (1999). In Sec. IV.3 a set of EOSs obtained within this last formalism in Malik et al. (2022) will also be introduced and referred to as DDB EOSs. In the above expression $\gamma^{\mu}$ and $\bm{t}$ designate, respectively, the Dirac matrices and the isospin operator. The vector meson field strength tensors are given by $F^{(\omega,\varrho)\mu\nu}=\partial^{\mu}A^{(\omega,\varrho)\nu}-\partial^{\nu}A^{(\omega,\varrho)\mu}$.
For the term $\mathcal{L}_{NL}$ we take
$\mathcal{L}_{NL}=$ $-\frac{1}{3}bg_{\sigma}^{3}(\sigma)^{3}-\frac{1}{4}cg_{\sigma}^{4}(\sigma)^{4}+\frac{\xi}{4!}(g_{\omega}^{2}\omega_{\mu}\omega^{\mu})^{2}$ (2)
$+\Lambda_{\omega}g_{\varrho}^{2}\bm{\varrho}_{\mu}\cdot\bm{\varrho}^{\mu}g_{\omega}^{2}\omega_{\mu}\omega^{\mu}.$
The couplings in front of each term, $b,\,c,$ $\xi$, $\Lambda_{\omega}$ are fixed by a given set of values together with the meson-nucleon couplings $g_{i}$ imposing the constraints defined in Table 1 within a Bayesian inference calculations, see Malik et al. (2023) for more details. This Lagrangian density is the starting point of many RMF models Horowitz and Piekarewicz (2001); Todd-Rutel and Piekarewicz (2005); Providência et al. (2014); Chen and Piekarewicz (2014); Bao and Shen (2015); Tolos et al. (2017); Fattoyev et al. (2020). The couplings in front of each term, $b,\,c,$ $\xi$, $\Lambda_{\omega}$ are fixed together with the meson-nucleon couplings $g_{i}$ imposing the constraints defined in Table 1 within a Bayesian inference calculations, see Malik et al. (2023) for more details. This Lagrangian density is the starting point of many RMF models Horowitz and Piekarewicz (2001); Todd-Rutel and Piekarewicz (2005); Providência et al. (2014); Chen and Piekarewicz (2014); Bao and Shen (2015); Tolos et al. (2017); Fattoyev et al. (2020).
The introduction of hyperons requires that the first term of the Lagrangian density, eq. (1), also sums over the hyperons. The couplings of the hyperons to the mesons were constrained by the properties of the hypernuclei and implemented as described in Malik and Providência (2022): for the vector isoscalar mesons, the SU(6) quark model values were chosen. For the vector isovector, the interaction is defined by the isospin projection multiplied by the $\varrho$ meson-nucleon coupling. The possible range of values for the coupling to the sigma meson has been obtained from a rather large number of RMF models by fitting some hypernuclei properties (see Fortin et al. (2017); Providência et al. (2018); Fortin et al. (2020)). Note that only the most probable hyperons, $\Lambda$ and $\Xi^{-}$, were considered. The experimental data seem to indicate that the $\Sigma$-nucleon interaction is repulsive, which prevents the onset of the $\Sigma^{-}$ hyperon inside NS.
After the detection of the two solar mass pulsar PSR J1614-2230 Demorest et al. (2010), it was argued that baryonic matter with hyperons would not support two solar mass NS, see discussion in Demorest et al. (2010); Vidana et al. (2011). However, it was later shown that two solar mass stars could also include hyperons not only within a RMF description considering the interaction constrained by hypernuclei properties Bednarek et al. (2012); Weissenborn et al. (2012); Providência et al. (2014); Oertel et al. (2015) but also within a microscopic non-relativistic description solving the Schrödinger equation by means of an auxiliary field diffusion Monte Carlo algorithm and including two-body and three body forces involving the $\Lambda$ hyperon Lonardoni et al. (2014). For a more complete overview of these subject, we refer the reader to the reviews Chatterjee and Vidaña (2016); Tolos and Fabbietti (2020); Burgio et al. (2021). In our study, we use datasets generated within the RMF description which allows for two solar mass stars with hyperons.
The data set, which is publicly available [^1] https://zenodo.org/records/7854112, contains 17810 EoS of pure nucleonic matter and 18728 EoS of hadronic matter containing both nucleons and hyperons.
The properties mass and radius of spherically symmetric and static relativistic stars in hydrostatic equilibrium are obtained by solving the Tolman-Oppenheimer-Volkov (TOV) equations Tolman (1939); Oppenheimer and Volkoff (1939)
$\frac{dP(r)}{dr}=\frac{\varepsilon(r)m(r)}{r^{2}}\left[1+\frac{P(r)}{\varepsilon(r)}\right]\\ \left[1+\frac{4\pi r^{3}P(r)}{m(r)}\right]\left[1-\frac{2m(r)}{r}\right]^{-1}\;,$ (3)
$\frac{dm(r)}{dr}=$ $4\pi r^{2}\varepsilon(r)\;.$ (4)
where $r$ is the radial distance from the NS center, $p(r)$ and $\varepsilon(r)$ are the pressure and energy density inside the NS at $r$, and $m(r)$ is the mass enclosed within the sphere of radius $r$. The tidal deformability, $\Lambda$, is expressed in terms of the dimensionless tidal Love number, $k_{2}$ , as follows
$\Lambda=\frac{2}{3}k_{2}(C)^{-5}\;,$ (5)
where $C=M/R$ is the compactness of the star. The second order Love number is then defined as
$k_{2}=$ $\frac{8C^{5}}{5}\left(1-2C\right)^{2}\left[2+2C\left(y_{R}-1\right)-y_{R}\right]\times$
$\{2C\left[6-3y_{R}+3C\left(5y_{R}-8\right)\right]$
$+4C^{3}\left[13-11y_{R}+C\left(3y_{R}-2\right)+2C^{2}\left(1+y_{R}\right)\right]$
$+3\left(1-2C\right)^{2}\left[2-y_{R}+2C\left(y_{R}-1\right)\right]\ln\left(1-2C\right)\}^{-1}\;,$ (6)
the boundary value of $y(r)$ at the surface, denoted as $y_{R}\equiv y(r=R)$ arises from the following first-order differential equation (Hinderer (2008))
$\frac{dy(r)}{dr}=-\frac{y(r)^{2}}{r}-\frac{y(r)}{r}F(r)-rQ(r)\;$ (7)
where $F(r)$ and $Q(r)$ being functions of $\varepsilon(r)$, $P(r)$, and $m(r)$, are represented as
$F(r)$ $=\left(1-4\pi r^{2}\left[\varepsilon(r)-P(r)\right]\right)\left[1-\frac{2m(r)}{r}\right]^{-1}\;,$ (8)
$Q(r)$ $=4\pi\left[5\varepsilon(r)+9P(r)+\frac{\varepsilon(r)+P(r)}{v_{s}^{2}(r)}-\frac{6}{r^{2}}\right]$
$\left[1-\frac{2m(r)}{r}\right]^{-1}-\frac{4m^{2}(r)}{r^{4}}\Biggl{[}1+\frac{4\pi r^{3}P(r)}{m(r)}\Biggr{]}^{2}$
$\left[1-\frac{2m(r)}{r}\right]^{-2}\;.$ (9)
These equations can be solved numerically simultaneously with the TOV Eqs. 3 and 4, also taking into account $y(r=0)=2$, as shown in Krastev (2022), along with the conditions for the center of the star. By doing so, the value of $y_{R}$ is determined and subsequently substituted into the Love number, Eq.II.1, to obtain the tidal deformability ($\Lambda$) using Eq.5.
In Fig.1, in the right plot, the $\beta$-equilibrium NS matter EOS with and without hyperons is shown. This underlying EOS, when integrated using the TOV equations alongside Eqs. 8 and 9, results in the mass-radius distribution present in the left plot and the corresponding mass-tidal deformability curves shown in the center plot.
To solve the TOV equations, a complete EOS with crust and core must be constructed. For the outer crust, the BPS EOS was considered. The inner crust was described by a polytropic function interpolating between the outer crust and the neutron star core at a density of 0.04fm-3, see Malik et al. (2023) for details. This prescription has been shown to give rise to radius uncertainties for low/medium mass stars below 200m, and a negligible effect for massive stars.
The inclusion of hyperons leads to a stiffening of the core EOS at low/intermediate densities, followed by a softening once hyperon onset occurs. The stiffening at low densities is essential for the model to reach $2M_{\odot}$ NSs. The behavior of the EOS at low and high densities is directly reflected in the mass-radius and mass-tidal deformability curves, in particular: i) low/medium-mass hyperonic stars can have larger radii and tidal deformabilities than nucleonic stars; ii) the inclusion of hyperons leads to lower maximum masses.
The Bayesian framework used to generate the two sets of EOSs imposes minimal constraints on various nuclear saturation properties. This approach ensures the reproduction of $2M_{\odot}$ NSs and maintains consistency with low-density pure neutron matter as predicted by N3LO calculations in chiral effective field theory (see table 1).
@@
Figure 1: Mass-radius, left plot, mass-log(tidal deformability), center plot, and pressure-energy density, right plot, curves for nucleonic matter (17810 curves, shown in blue) and hyperonic matter (18728 curves, shown in red).
@@ IV.4 Application to real data
To the best of our knowledge, our work represents the first attempt to frame the detection of hyperons as a classification problem. However, there have been attempts to identify the characteristics which define the properties of NSs that contain hyperons in their composition. For instance, in the study Fortin et al. (2015), where the authors study 14 theoretical EOS of dense matter based on the RMF framework, and analyse the effects of the presence of hyperons, they conclude that medium mass stars with hyperons have larger radii than pure nucleonic stars with the same mass. This is also a feature of the datasets we have used to define our classification models. In another study, Sun et al. (2023b), the authors study the implications of considering the presence of hyperons within NS on the hyperon couplings. They apply a Bayesian inference calculation to constrain the hyperon couplings from hypernuclear properties and astrophysical observations of NS, and conclude that the former are much more constraining than the astrophysical data. This seems to indicate that in order to identify hyperons inside NS measurements with smaller uncertainties are needed. Another potential indicator of hyperon presence in NSs comes from the cooling curves. In a model that does not predict nucleonic direct Urca, because the proton fraction remains too small in the NS interior, the opening of direct Urca processes may occur only after the hyperon onset, allowing rapid cooling in the NS Providência et al. (2018); Fortin et al. (2020). However, the cooling curves are subject to much larger uncertainties than the NS radii, and therefore the information we can extract from the cooling curves is still limited.
@@ V Conclusions
We investigated a classification problem aimed at determining the potential presence of hyperons within NSs based on their macroscopic properties. We investigate the capacity of NNs for the present binary classification problem of detecting the presence of hyperons. The dataset used has been constructed from an RMF approach within a Bayesian framework, incorporating crucial constraints derived from both nuclear matter properties and NS observations. The dataset consists of 17810 EOSs for pure nucleonic matter and 18728 EOSs including hyperons. This approach was chosen to achieve our primary objective: to investigate the presence of hyperons within NSs based on observational data, which requires the use of microscopic models. From this comprehensive dataset, we generated three distinct sets by selecting either 5, 10, or 15 pairs of mass-radius ($\bm{R}$), mass-tidal deformability ($\bm{\Lambda}$), or both ($\bm{R\Lambda}$) as input parameters. We varied the input noise to create three different datasets for each input size, as summarized in Table II. These sets represent different levels of noise scatter in the input data. Consequently, a total of 27 NN models were trained, each corresponding to one of these possible combinations. We conducted a grid search to optimize the model architecture specifically for the $R_{0}$, $\Lambda_{0}$, and $R\Lambda_{0}$ sets. Once the optimal architecture was identified, we evaluated the model’s performance using confusion matrices and key metrics such as accuracy and F1 score. To ensure that the model’s performance is not biased by a fixed threshold, we also calculated the AUC for accuracy as a function of the threshold. The AUC values demonstrated consistent behavior with the accuracy measured at the standard threshold of $0.5$ Our findings reveal that as input noise increases, the model requires a larger input size to maintain high accuracy. In contrast, for datasets without noise, the model quickly reaches peak accuracy with a smaller number of input pairs. This suggests that for observational data—roughly equivalent to the sets with $\mathbb{X}_{2}$ noise levels—larger input sizes are needed to achieve accuracy comparable to the noise-free case. Nevertheless, our model achieves high scores across all evaluated metrics, including in scenarios with the highest input noise. We further tested the model on a different dataset, also generated using an RMF approach but with density-dependent couplings and a different particle composition. The results demonstrated good accuracy across all models, consistently exceeding 73%. For the AUC value and the F1 score, the performance decreased slightly. The F1 score is particularly significant given the highly imbalanced nature of this test set, and has indicated that the model is effectively handling and correctly predicting outcomes despite the data imbalance. While the model performs well across most datasets, there is room for improvement, particularly in handling the $\Lambda$ datasets with higher noise levels. Finally, we tested our models on real observational data for the $\bm{R}$ models with $Q=5$, using a subset of observations that lie within the model’s training region. The results revealed that the majority of predictions fell within the region without hyperons, suggesting that hyperons may be absent inside NSs. More specifically, 65% of the data across all datasets indicated predictions in the region without hyperons. We also conducted tests with fewer samples for the $R_{2}$ set at $Q=10$ and $Q=15$, achieving solid results that consistently indicate the absence of hyperons. The interpretability of machine learning models is an area of active research. In future work, we plan to conduct feature importance analyses on our current model to identify the regions of the $M(R)$ diagram that are most informative for detecting hyperonic degrees of freedom.
@@
Looking ahead, upcoming observations will enable us to test models that incorporate tidal deformability, potentially providing additional insights into NS compositions. A significant direction for future work involves the application of deep ensembles or the incorporation of BNNs. These approaches would allow us to quantify both epistemic and aleatoric uncertainties, providing a more comprehensive understanding of model reliability. Given the results obtained with different nuclear models it could also be advantageous to perform a more rigorous model optimization, such as using Bayesian search methods, to enhance performance and accuracy. However, this approach can be resource-intensive in terms of both time and memory. Additionally, expanding the model to consider the presence of a quark phase in NS compositions and training and testing it under these new conditions could offer intriguing insights into the physics of dense matter. The codes to reproduce the present study are available on request.